OpenStudy (anonymous):

Help? An Eskimo is standing on a floating iceberg headed due east at 8 mph. If the stream is flowing due north at 5 mph, find the velocity (magnitude) of the iceberg. Show all work

OpenStudy (anonymous):

The speeds represent the two legs of a triangle. Using the Pythagorean Theorem we get 64+25= x^2...solving for x yields an answer of approximately 10.44 mph northeast.

OpenStudy (anonymous):

It seems to have cut off the full question...

OpenStudy (anonymous):

An Eskimo is standing on a floating iceberg headed due east at 8 mph. If the stream is flowing due north at 5 mph, find the resultant direction and velocity (magnitude) of the iceberg. Show all work.

OpenStudy (anonymous):

Resultant Direction = √ ( 8^2 + 5^2) Resultant Direction = √(64+25) Resultant Direction = √89 Resultant Direction = 9.43 mph is what i have so far

OpenStudy (anonymous):

help?

OpenStudy (anonymous):

okay i will

OpenStudy (anonymous):

Well, this is difficult to explain without drawing, but suppose You know how to draw 2D vectors. One arrow pointing right of length 8 (representing movement due east) and another arrow pointing upwards of length 5 (representing movement due north). Both arrows start from the same initial (reference) point, representing the iceberg. Their resultant is a diagonal arrow pointing northeast. Its length is sqrt(5^2+8^2) = 9.434, so this is the resultant speed (velocity magnitude) in mph. As to the direction, the angle between the north and the resultant determines its direction: 5/8 = tan(90deg-angle) This should be obvious from the drawing. Now: arctan(5/8) = 90deg-angle from which the angle can be calculated to be 58 degrees. Hence, the answer, if I didn't make any calc mistakes, should be exactly 58 degrees northeast. The role of the eskimo there is not very clear (?)

OpenStudy (anonymous):

haha okay thank you so much you really helped me out!!!